OpenStudy (anonymous):

3. how many solution does the system of equations have? x+-4y+4 and 2x+8y+8 4. how many solutions does system of equations have? y=6x+2 and 5x-20y=60 5. how many solutions does system of equations have? x-4y=12 and 5x-20y=60 6. how many solutions does system of equations have? y-7x=-14 and 7y-49x+-2

OpenStudy (anonymous):

14. what is the solution of the system? 2x-y+-7 4x-y=-4 A. (-1.5,4) B. (1.5,10) C. (4,-1.5) D. (-1.5,2) 15.what is the solution of the system? 5x+6y=20 9x+11y=32 A.(28,-20) B.(28,20) C.(-28,20) D.(20,-28) if your willing to help me on more I will medal and fan but I need only right answers

OpenStudy (anonymous):

@CloverRacer @SolomonZelman @Hero @parthkohil

OpenStudy (anonymous):

if you help me now I give medal to anybody that is sure that they know the right anwsers